OpenStudy (like1234):

What property is shown? –12 + (27 + 12) = –12 + (12 + 27) A. associative property B. commutative property C. distributive property D. opposite of a sum property

OpenStudy (like1234):

Will medal!!!!!!!!!!

OpenStudy (like1234):

I think B

OpenStudy (anonymous):

The answer would be commutative because it changes order.

OpenStudy (like1234):

ok thx!

OpenStudy (like1234):

so is the answer for this on B Which is an example of the commutative property? A. –13 + 27 = 27 + (–13) B. –21 + (11 + 13) = (–21 + 11) + 13 C. 4(–3 + 5) = 4(–3) + 4 ∙ 5 D. –(–8 + 3) = 8 + (–3)

OpenStudy (mathstudent55):

commutative: change of order associative: change of grouping

OpenStudy (like1234):

ok thx!

OpenStudy (mathstudent55):

distributive: a(b + c) = ab + ac

OpenStudy (anonymous):

Associative property keeps the order but changes the parentheses.

OpenStudy (like1234):

ok...................

OpenStudy (like1234):

so am i correct???

OpenStudy (mathstudent55):



OpenStudy (anonymous):

No its not B. B is associative property

OpenStudy (like1234):

Oh

OpenStudy (mathstudent55):

Commutative property is a change of order. Here is the general way of stating the commutative property of addition: a + b = b + a

OpenStudy (like1234):

wit what problem is this for?

OpenStudy (like1234):

*wait

OpenStudy (anonymous):

I am talking about the second problem

OpenStudy (like1234):

ok

OpenStudy (mathstudent55):

For the second problem, in B., the 11 and 13 were grouped together on the left side. On the right side, the -21 and 11 are grouped together. This is a change of grouping, so B is associative, not commutative.

OpenStudy (like1234):

so for communicative u have 2 change the whole thing?

OpenStudy (mathstudent55):

You need to find a statement that is an example of a + b = b + a, which is commutative.

OpenStudy (like1234):

OH

OpenStudy (mathstudent55):

It is not "communicative." Communicative means you like to communicate. It is "commutative." Commutative means you change, as in change the order.

OpenStudy (like1234):

ok im confused for communicative u change the order the opposite

OpenStudy (like1234):

wit my problem is commutative property not communicative

OpenStudy (like1234):

*wait

OpenStudy (mathstudent55):

Commutative: See if one of the choices shows this: (first number) + (second number) = (second number) + (first number)

OpenStudy (like1234):

oh i see thx

OpenStudy (like1234):

so A then right?

OpenStudy (mathstudent55):

Commutative property of addition. (first number) + (second number) = (second number) + (first number) a + b = b + a Example: 3 + 5 = 5 + 3

OpenStudy (mathstudent55):

Correct. A is the answer.

OpenStudy (like1234):

k

OpenStudy (like1234):

thx sooooo much

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (like1234):

What value of q makes the statement true? –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17

OpenStudy (mathstudent55):

yw

OpenStudy (like1234):

i now it has 2 be 3 but im not sure if it is -3 or 3

OpenStudy (like1234):

@mathstudent55

OpenStudy (like1234):

@KAKES1967

OpenStudy (like1234):

I think just 3

OpenStudy (mathstudent55):

Let's use the distributive property first on the left side. –(q + 14) = –3 + (–14) -q + (-14) = -3 + (-14)

OpenStudy (mathstudent55):

Now you see that -q must equal -3 That means q = 3, not -3

OpenStudy (mathstudent55):

You are correct. q = 3

OpenStudy (anonymous):

Yes you are correct but let @mathstudent55 explain to you why its that

OpenStudy (like1234):

thx guys!

OpenStudy (mathstudent55):

yw

OpenStudy (like1234):

ok last thing Use an addition property to solve for b. –12 + 16 = 16 + b A. –28 B. –12 C. 4 D. 12

OpenStudy (like1234):

i think B

OpenStudy (anonymous):

Correct

OpenStudy (like1234):

thx

OpenStudy (mathstudent55):

You want b alone on the right side. Since 16 is being added to b, you need to add something to 16 to make it zero. Additive inverses add to zero.

OpenStudy (anonymous):

Its using communative property

OpenStudy (like1234):

Got 100% thx sooooooooo much u guys really helped

OpenStudy (anonymous):

Your welcome :P

OpenStudy (anonymous):

Good job

OpenStudy (mathstudent55):

If you add -16 to 16 you get zero because -16 is the additive inverse of 16. The rule with an equation is you must add the same to both sides of the equation. Since you need to add -16 to the right side, you must add -16 to the left side too.

OpenStudy (like1234):

ok

OpenStudy (mathstudent55):